OpenStudy (anonymous):

6= 2(x+8)-5x = 3 1/3 am i right?

OpenStudy (anonymous):

right!! well done!

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

6= 2(x+8)-5x 6= 2x+16-5x 6 = 16 - 3x -16 -16 -10 = -3x -10/-3 = x x = 10/3 = 3 1/3 you got an A

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

thankss